OpenStudy (anonymous):

Solve by completing the square. 5x^2-2x+1=0 i just can't figure this out.

OpenStudy (anonymous):

you have to bring it to theform (x - a)^2 = b^2

OpenStudy (anonymous):

\[x ^{2}-\frac{ 2 }{5 }x+\frac{ 1 }{ 5 }=0\] \[x ^{2}-\frac{ 2 }{5 }x+\left( \frac{ -2 }{ 5*2 } \right)^{2}-\left( \frac{ -2 }{ 5*2 } \right)^{2} +\frac{ 1 }{ 5} =0\] \[\left( x-\frac{ 1 }{ 5 } \right)^{2}-\frac{ 1 }{ 25 }+\frac{ 1 }{ 5 }=0\] \[\left( x-\frac{ 1 }{5 } \right)^{2}+\frac{ -1+5 }{ 25 }=0\] \[\left( x-\frac{ 1 }{5 } \right)^{2}-\frac{- 4 }{ 25 }=0\] \[\left( x-\frac{ 1 }{ 5 } \right)^{2}-\frac{ 4\iota ^{2} }{25 }=0\] \[\left( x-\frac{ 1 }{5 } \right)^{2}-\left( \frac{ 2\iota }{ 5 } \right)^{2}=0\] \[\left( x-\frac{ 1 }{ 5 }+ \frac{ 2\iota }{ 5 }\right)\left( x-\frac{ 1 }{ 5 }-\frac{ 2\iota }{ 5 } \right)=0\]

OpenStudy (anonymous):

super factoring :)